# DeepVoting: Learning Voting Rules with Tailored Embeddings

### v1

@@ Abstract
Aggregating the preferences of multiple agents into a collective decision is a common step in many important problems across areas of computer science including information retrieval, reinforcement learning, and recommender systems. As Social Choice Theory has shown, the problem of designing algorithms for aggregation rules with specific properties (axioms) can be difficult, or provably impossible in some cases. Instead of designing algorithms by hand, one can learn aggregation rules, particularly voting rules, from data. However, the prior work in this area has required extremely large models, or been limited by the choice of preference representation, i.e., embedding. We recast the problem of designing a good voting rule into one of learning probabilistic versions of voting rules that output distributions over a set of candidates. Specifically, we use neural networks to learn probabilistic social choice functions from the literature. We show that embeddings of preference profiles derived from the social choice literature allows us to learn existing voting rules more efficiently and scale to larger populations of voters more easily than other work if the embedding is tailored to the learning objective. Moreover, we show that rules learned using embeddings can be tweaked to create novel voting rules with improved axiomatic properties. Namely, we show that existing voting rules require only minor modification to combat a probabilistic version of the No Show Paradox.
@@ 1 Introduction
Research in Computational Social Choice (COMSOC) and Algorithmic Game Theory (AGT) focuses heavily on the design and analysis of mechanisms for collective decision making. Canonically, agents arrive with individual preferences over a set of alternatives or outcomes, and a centralized mechanism must aggregate these preferences into a shared choice (voting and selection) or allocation of items (matching and auctions) Shoham and Leyton-Brown (2008). The goal is to design mechanisms with certain desirable properties, characterized by axioms; i.e. optimizing a particular objective or satisfying certain constraints.
@@
One of the central results in social choice is Arrow’s General Impossibility Theorem Arrow et al. (1963), which identifies a set axioms that no collective mechanism can satisfy at once. Following Arrow, decades of research has produced myriad theorems showing which axioms are satisfied by which mechanisms and which sets of axioms lead to an impossibility result Sen (2018). On the algorithmic and computational side this includes properties of optimality and computational complexity Brandt et al. (2016).
@@
Finding rules that satisfy a given set of axioms can be difficult, especially when it is unknown if such a rule exists. Hence, recent work has turned to machine learning techniques to design novel mechanisms. This idea has been applied in areas from auctions to voting rules to matchings Xia (2013); Sandholm (2003); Curry et al. (2022); Ravindranath et al. (2021). Previous work on learning voting rules has been hampered by technical challenges, including extremely large/sophisticated neural nets Anil and Bao (2021), limited data Burka et al. (2022), or failure to account for the full consequences of the design choices Firebanks-Quevedo (2020).
@@
Our approach to improving the learning of existing and novel voting rules is to use hand-crafted embeddings derived from the social choice literature. As we show, these embeddings enable fast learning with fewer parameters and the ability to scale to larger populations of voters. Our embeddings reduce the input size of our neural net, greatly reducing the number of model parameters. Anil and Bao (2021) found that using a multi-layer perceptron (MLP), i.e., a deep neural network, to learn voting rules was limited by the network’s fixed input size, and so they introduced more sophisticated architectures to permit scaling and pad smaller profiles. In their tests to see which architectures scaled best with the number of voters, their MLP architecture was necessarily excluded. By contrast, our embeddings enable the size of the input layer of the MLP to be independent of the number of voters, thereby enabling greater generalization with scale. Similarly, when we train or test our model with preference profiles containing fewer voters, our input does not require any form of padding. As we show, when using embeddings to learn particular rules or axioms, the choice of embedding must correspond to the learning objective.
@@
Sen (2018) observed that mechanisms can be examined in terms of what information they use and ignore from preference profiles. A key contribution of our work is a more complete understanding of the relationship between the choice of embedding and the resulting learnability and efficiency of the mechanisms. We observe that, like any compression algorithm, embeddings can be lossy and impose a bound on the learnability of rules and axioms. Indeed, we show that certain embeddings lose the required information needed to learn particular rules. Our experiments have also lead us to pose new theoretical questions about the information preservation of embeddings.
@@
When most people think of voting they think of classical deterministic rules that take in preferences over a small set of candidates and return a single winner Zwicker (2016); Taylor (2005). However, the full space of social choice mechanisms is much richer with mechanisms varying by the data types of their inputs and outputs. For example, voters may give approval ballots, rankings, or weightings to different candidates; the outcome of the mechanism may be a single winning candidate, collection of winners, or ordering of the candidates, among other options.
@@
We study probabilistic social choice functions (PSCFs) which take a profile over candidates and return a lottery (probability distribution) over the candidates Brandt (2017). PSCFs offer several advantages when learning voting rules with neural networks as compared to single-winner voting rules. First, since we are outputting a lottery over the alternatives, we are not as concerned with tie-breaking. Tie-breaking introduces complications into the design and analysis of voting rules Aziz et al. (2013b), and random tie-breaking is not learnable. Second, PSCFs provide a natural connection between the discrete formulations of rules and axioms and the construction of continuous loss functions for training based on divergences between distributions, e.g. L1.
@@
After first showing that existing rules can be efficiently learned using tailored embeddings, we address the challenge of taking existing rules and tweaking them to improve their axiomatic properties. In particular, we focus on the No Show Paradox in which a voter can induce an outcome they prefer by abstaining instead of voting Moulin (1988). Certain single-winner rules, like Plurality, Borda, and Simpson-Kramer are known to satisfy the Participation axiom, so no voter can be made better off by abstaining and the paradox does not arise. However, most other common voting rules exhibit this paradox. In our experiments, we take models trained to learn existing PSCFs and retrain them using a loss function that adds in a continuous relaxation of the Participation axiom and show which rules can be adjusted to be more resistant to the paradox without sacrificing accuracy.
@@
We choose the Participation axiom in particular because it is an inter-profile axiom, which requires reasoning about counterfactuals on what the preference profile could have otherwise been had the voters behaved differently. Inter-profile axioms are particularly challenging for learning from data because they increase the computational complexity of computing loss functions. It also requires that the model be able to take profiles of different sizes (differing by one voter) as input, which our embeddings enable us to do. Since abstention can be a strategic behavior by voters, learning the Participation axiom is closely related to Automated Mechanism Design, which focuses on creating desirable economic mechanisms for strategic agents and “shifts the burden of design from man to machine” Sandholm (2003).
@@ 2 Related Work
Xia (2013) and Procaccia et al. (2009) proposed incorporating voting axioms into a machine-learning framework as a means of evaluating learned social choice mechanisms. In the space of auction design and matching there has been work on using neural nets for better mechanisms Dütting et al. (2019); Pavlov (2011); Malakhov and Vohra (2008) including learning new types of auction mechanisms Curry et al. (2022) as well as complex preference structures Peri et al. (2021). More recently, the work of Ravindranath et al. (2021) has looked at how to learn new allocation mechanisms that bridge the gap between stability (as compared to the deferred acceptance algorithm Gale and Shapley (1962)) and strategyproofness (as compared to random serial dictatorship (RSD) Aziz et al. (2013a)). While the work of Ravindranath et al. (2021), Firebanks-Quevedo (2020), and most recently Anil and Bao (2021) has shown promise for learning mechanisms, these efforts do not closely consider the role of embeddings.
@@
While formal proposals to learn voting rules date back over a decade Xia (2013), considerable attention to learning voting rules has increased in recent years. Kujawska et al. (2020) and Burka et al. (2022) used several common machine learning methods to mimic existing voting rules. However, both of these works overlooked the importance of the choice of embedding in the role of learning, finding that certain rules were “easier” to learn. Subsequently, Anil and Bao (2021) showed that PIN architectures have the advantage of better generalization to larger numbers of voters. We build on this work by showing that we can achieve high accuracy efficiently with small MLPs by using hand-crafted embeddings.
@@
Procaccia et al. (2009) showed that positional scoring rules are efficiently PAC learnable, but learning pairwise comparison-based voting rules requires an exponential number of samples. While we do not escape the asymptotic limits, we examine two embeddings based on tournament graphs that are design to facilitate more efficient learning of pairwise-comparison based rules. Firebanks-Quevedo (2020) use one measure of optimality (Condorcet consistency) and strategyproofness for learning. However, the manipulations of the chosen embedding cannot learn strategyproofness, leading to poor results. Finally, Wilson (2019) focus on the problem of learning a voting rule given a set of pair-wise relations and properties that must hold for the optimization criteria. However, this work is focused on the possibility of learning these functions and does not employ any machine learning techniques.
@@
The loss function chosen by Armstrong and Larson (2019) was a function of the profile and outcome, and thus could learn a rule but not inter-profile axioms like Participation. Recently, Mohsin et al. (2022) focused on the problem of designing and/or learning fair and private rules, proving that under measures of differential privacy there is an upper bound on the trade-off between group fairness and efficiency.
Learning voting rules bears some similarity to the well studied area of learning to rank (L2R) from the machine learning literature Cao et al. (2007). In L2R one typically is concerned only with accurate recovery of the population preference and not the axioms or properties of the aggregation method itself (e.g., fairness). Indeed, one can think of our work enforcing inter-profile axioms on the learned aggregation procedures as an important step.
@@ 3 Preliminaries
Let $V$ be a set of $n$ voters and $C$ a set of $m$ candidates. Each voter $i\in V$ reports a strict order $x_{i}$ over all the candidates in $C$ as their ballot. We will denote that $i$ strictly prefers $a$ over $b$ by $a\succ_{i}b$ for $a,b\in C$. There are $m!$ possible ballots, or ways to strictly order (permute) the candidates in $C$. A list of $n$ ballots, one for each voter, constitutes a profile $X=(x_{i})_{i\in V}$. Voter $i$ ranks candidate $a$ at position $x_{i}(a)\in[m]$, using $[k]=\{1,\ldots,k\}$. Let $\mathcal{X}$ be the set of all possible profiles.
@@
A probabilistic social choice function (PSCF) is a function $f:\mathcal{X}\rightarrow\Delta(C)$ that takes a profile $X\in\mathcal{X}$ as input and returns a lottery, or probability distribution $f(X)\in\Delta(C)$ over the set of candidates in the profile, where $\Delta(C)$ is the set of all lotteries over $C$. Let $\mathcal{F}$ be the set of all such PSCFs.
Any PSCF can be used to construct a non-deterministic voting rule by sampling a winner from the lottery. A PSCF that places all probability mass on a single candidate for all profiles is deterministic. If the candidate that receives all of the probability mass is the same candidate for all profiles, then it is a dictatorial voting rule. Many PSCFs we consider return for all profiles a lottery that is a uniform distribution over a non-empty subset of the candidates, i.e., there are multiple potential winners that we would have to choose from for a single-winner voting rule. Therefore, let $U(Y)$ denote the uniform distribution over any finite set $Y$. When referring to lotteries over candidates, we let $U(Y)$ denote the distribution that is uniform over $Y\subseteq C$ and zero on $C\backslash Y$.
Embedding
Simple feed-forward neural networks require a fixed-size input for learning and inference, corresponding to the size of their input layer. If we were to learn voting rules using neural networks that take the entire profile as input, then not only does the input layer need to be large $(m\times n)$, but it also prevents scaling up as the number of voters grows. Similarly, if the number of voters shrinks, then the profile would have to be padded carefully in a way that doesn’t negatively impact the model. If we want to learn rules that are agnostic to the number of voters, we need to construct embeddings of fixed-size that retain the relevant information for profiles with any number of voters. Naturally, different embeddings preserve different information from the original profile, leading them to different efficacy when learning different rules and axioms. Note that most rules and axioms in the literature are defined for any positive number of voters, so we would like our learned mechanisms to be similarly agnostic.
Figure 1: Each of our three embeddings derived from a ballot profile. Note that the size of the profile (a) grows in $O(mn)$, while each of our embeddings grows with $O(m^{2})$, which is far smaller when $m<<n$. However, our embeddings do not always preserve all of the information in the original profile.
An embedding $T$ is a function $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ mapping profiles to some codomain $\mathcal{X}^{\prime}$. The embeddings we are concerned with are many-to-one mappings. This means multiple different profiles may have the same embedding, i.e. $T(X)=T(\tilde{X})$ for some $X,\tilde{X}\in\mathcal{X}$ where $X\neq\tilde{X}$. In other words, $T$ will not be reversible, and $T(X)$ will not preserve all information about $X$. We denote by $\mathcal{F}^{\prime}$ the set of all probabilistic functions of the form $f^{\prime}:\mathcal{X}^{\prime}\rightarrow\Delta(C)$. Note that while we designate $\mathcal{X}$ to always contain strict orders over candidates, the structure of $\mathcal{X}^{\prime}$ will be different for different embeddings. Our three embeddings are drawn from the voting literature, but are not commonly recognized as embeddings in the machine learning literature.

### v2

# DeepVoting: Learning and Fine-Tuning Voting Rules with Canonical Embeddings

@@ Abstract
Aggregating agent preferences into a collective decision is an important step in many problems (e.g., hiring, elections, peer review) and across areas of computer science (e.g., reinforcement learning, recommender systems). As Social Choice Theory has shown, the problem of designing aggregation rules with specific sets of properties (axioms) can be difficult, or provably impossible in some cases. Instead of designing algorithms by hand, one can learn aggregation rules, particularly voting rules, from data. However, prior work in this area has required extremely large models or been limited by the choice of preference representation, i.e., embedding. We recast the problem of designing voting rules with desirable properties into one of learning probabilistic functions that output distributions over a set of candidates. Specifically, we use neural networks to learn probabilistic social choice functions. Using standard embeddings from the social choice literature we show that preference profile encoding has significant impact on the efficiency and ability of neural networks to learn rules, allowing us to learn rules faster and with smaller networks than previous work. Moreover, we show that our learned rules can be fine-tuned using axiomatic properties to create novel voting rules and make them resistant to specific types of “attack”. Namely, we fine-tune rules to resist a probabilistic version of the No Show Paradox.
@@ 1 Introduction
Computational Social Choice (COMSOC) and Algorithmic Game Theory (AGT) focus heavily on the design and analysis of mechanisms for collective decision making. Canonically, agents arrive with individual preferences over a set of alternatives or outcomes, and a mechanism aggregates these preferences into a shared choice (voting and selection) or allocation (matching and auctions) (Shoham and Leyton-Brown 2008). The goal is to design mechanisms with certain desirable properties, characterized by axioms; i.e. optimizing a particular objective or satisfying certain constraints.
@@
A central result in Social Choice is Arrow’s General Impossibility Theorem (Arrow 1963), which identifies a set of axioms that no collective choice mechanism can satisfy simultaneously. Following Arrow, decades of research has produced myriad theorems showing which axioms are satisfied by which mechanisms and which lead to an impossibility results (Sen 2018), including optimality, computational complexity, and strategyproofness (Brandt et al. 2016).
@@
Finding rules that satisfy a given set of axioms can be difficult, especially when it is unknown if such a rule exists. Hence, recent work has turned to machine learning techniques to design novel mechanisms. This idea has been applied to auctions, voting rules, matchings, and beyond (Xia 2013; Sandholm 2003; Curry et al. 2022; Ravindranath et al. 2021). Previous work on learning voting rules has been hampered by technical challenges, including extremely large/sophisticated neural nets (Anil and Bao 2021), limited data (Burka et al. 2022), or failure to account for the full consequences of the design choices (Firebanks-Quevedo 2020).
@@
We improve the learning of existing and novel voting rules using common embeddings from the social choice literature. These embeddings enable faster learning with fewer parameters and scale to larger voter populations with better accuracy. Anil and Bao (2021) observed that using a multi-layer perceptron (MLP), i.e., a neural network, to learn voting rules was hampered by the network’s fixed input size, thus requiring more sophisticated architectures to permit scaling. Our embeddings reduce the input size of our neural net, greatly reducing the number of model parameters.
@@
Sen (2018) observed that aggregation mechanisms can be described by what information they use or ignore from preference profiles. A key challenge in designing neural networks for voting rules is understanding how to handle different numbers of voters or candidates, and several embeddings proposed in the social choice literature may provide a solution. In particular, embeddings whose size is independent of the number of voters can improve learning, but the choice of embedding must correspond to the learning objective. Like any compression algorithm, embeddings can be lossy and impose a bound on the learnability of rules and axioms. A key contribution of our work is a more complete understanding of the relationship between the choice of embedding and the resulting learnability and efficiency of the mechanisms. Our experiments inspire new theoretical questions about the information preservation and choice of embeddings.
@@
Often, when people think of voting they think of classical deterministic rules that take in preferences over a small set of candidates and return a single winner (Zwicker 2016; Taylor 2005). However, the full space of social choice mechanisms is much richer with mechanisms varying by the data types of their inputs and outputs; voters may give approval ballots, rankings, scores, or weightings to different candidates; the outcome of the mechanism may be a single winning candidate, collection of winners, or ordering of the candidates.
@@
We study probabilistic social choice functions (PSCFs) which take a profile over candidates and return a lottery (probability distribution) over the candidates (Brandt 2017). Unlike single-winner rules, PSCFs provide a natural connection between the discrete nature of rules and axioms and the continuous loss functions for training based on divergences between distributions. We use the L1 loss, the sum of point-wise absolute differences (taxicab distance) between the source (neural network result) and target distribution (voting rule) (Abu-Mostafa, Magdon-Ismail, and Lin 2012).
@@
We explore how well we can learn standard voting rules with common embeddings, testing against profiles both in and out of distribution. We then address the challenge of fine-tuning these networks to improve their axiomatic properties. We focus on the No Show Paradox in which a voter can induce an outcome they prefer by not voting (Moulin 1988). Single-winner Plurality, Borda, and Simpson-Kramer are known to satisfy this Participation axiom, though many other common voting rules are vulnerable to it (Zwicker 2016). We take models for PSCFs and fine-tune them using a loss function that adds in a continuous relaxation of the Participation axiom, showing rules can be refined to be more resistant to the paradox and maintain accuracy.
@@
We choose the Participation axiom because it is an inter-profile axiom, which requires reasoning about counter-factuals on what the preference profile could have been had the voters behaved differently. Inter-profile axioms are particularly challenging for learning from data as we must consider many different profiles, e.g., all $m!$ manipulations, in training. It also requires that the model be able to take profiles of different sizes (differing by one voter) as input, which our embeddings enable us to do. Since abstention can be a strategic behavior by voters, our work is closely related to Automated Mechanism Design, which aims to create desirable mechanisms for strategic agents and “shifts the burden of design from man to machine” (Sandholm 2003).
Critiques of using machine learning methods for voting rules include (1) most voting rules are simple to compute, why complicate it? and (2) how do we explain these rules if they are the output of a network? In regards to (1) we take an engineering approach: the first part of this paper is a study on how effectively we can learn these rules, so that we can then judge how well our more participation-proof rules work. In answer to (1) and (2) we agree that for many cases a direct implementation of the rule may be better. However, in some cases like recommender systems (Aird et al. 2024; Patro et al. 2020), where we want to optimize an objective, and limit our downsides, one may be okay with using a less explainable rule. [^1] Note that run-time efficiency is a key metric for recommender systems, and inference of our models is extremely fast. Ultimately, we want to learn novel rules that sit at the empirical Pareto front of an optimization criteria (e.g., top-cycle or Condorcet consistency) and resistance to forms of attack (manipulation, strategic abstention). This work is a concrete step in that direction, showing the limits of learning, and pointing out ways forward.
Contribution.
We (1) explicitly characterize which common embeddings from Social Choice are able to retain desirable properties and can be used to learn popular voting rules; (2) demonstrate, for the first time, that standard embeddings from Social Choice dramatically reduce the complexity and increase the efficiency of learning voting rules; (3) use transfer learning (fine-tuning) to add an axiomatic property to a learned voting rule, thereby making existing rules more resistant to strategic manipulation; and (4) provide strong evidence that training on Impartial Culture preferences teaches rules to generalize to additional preference distributions.
@@ 2 Related Work
Xia (2013) and Procaccia et al. (2009) proposed incorporating voting axioms into a machine-learning framework as a means of evaluating learned social choice mechanisms. In the space of auction design and matching there has been work on using neural nets for better mechanisms (Dütting et al. 2019; Pavlov 2011; Malakhov and Vohra 2008) including learning new types of auction mechanisms (Curry et al. 2022) as well as complex preference structures (Peri et al. 2021). More recently, the work of Ravindranath et al. (2021) has looked at how to learn new allocation mechanisms that bridge the gap between stability (as compared to the deferred acceptance algorithm (Gale and Shapley 1962)) and strategyproofness (as compared to random serial dictatorship (RSD) (Aziz, Brandt, and Brill 2013)). While the work of Ravindranath et al. (2021), Firebanks-Quevedo (2020), and most recently Anil and Bao (2021), has shown promise for learning mechanisms, these efforts do not closely consider the role of embeddings. Armstrong (2025) considered the impact of embeddings on rule learnability across a wide range of preference distributions but used a very limited network size and provided little subsequent analysis.
@@
While formal proposals to learn voting rules date back over a decade (Xia 2013), considerable attention to learning voting rules has increased in recent years. Kujawska, Slavkovik, and Rückmann (2020) and Burka et al. (2022) used several common machine learning methods to mimic existing voting rules. However, both of these works overlooked the importance of the choice of embedding in the role of learning, finding that certain rules were “easier” to learn but not theoretically characterizing why certain embeddings maintain properties, as we do. Subsequently, Anil and Bao (2021) showed that PIN architectures offer better generalization to larger numbers of voters. We build on this work by showing that we can achieve high accuracy efficiently with smaller MLPs by using specific embeddings.
@@
Procaccia et al. (2009) showed that positional scoring rules are efficiently PAC learnable, but learning pairwise comparison-based voting rules requires an exponential number of samples. While we do not escape the asymptotic limits, we examine two embeddings based on tournament graphs that facilitate more efficient learning of pairwise-comparison based rules. Firebanks-Quevedo (2020) uses one measure of optimality (Condorcet consistency) and strategyproofness for learning. However, as we show, the chosen embedding in that work cannot learn strategyproofness, leading to poor results. Finally, Wilson (2019) focus on learning a voting rule given pair-wise relations and properties that must hold for the optimization criteria. However, they focused on the possibility of learning these functions and does not employ any ML techniques.
@@
The loss function chosen by Armstrong and Larson (2019) was a function of the profile and outcome, and thus could learn a rule but not inter-profile axioms such as Participation. Recently, Mohsin et al. (2022) focused on the problem of designing and/or learning fair and private rules using random forests and a subset of embeddings we study, proving that under differential privacy there is an upper bound on the trade-off between group fairness and efficiency. Learning voting rules bears some similarity to the well studied area of learning to rank (L2R) from the machine learning literature (Cao et al. 2007). L2R is concerned with accurate recovery of the population preference and not the axioms or properties of the aggregation method itself (e.g., fairness). Indeed, one can think of our work enforcing inter-profile axioms on the learned aggregation procedures as an important step.
@@ 3 Preliminaries
Let $V$ be a set of $n$ voters and $C$ a set of $m$ candidates. Each voter $i\in V$ reports a strict order $x_{i}$ over all candidates in $C$ as their ballot. We denote that $i$ strictly prefers $a$ over $b$ by $a\succ_{i}b$ for $a,b\in C$. There are $m!$ possible ballots, or ways to strictly order (permute) the candidates in $C$. A list of $n$ ballots, one for each voter, constitutes a profile $X=(x_{i})_{i\in V}$. Voter $i$ ranks candidate $a$ at position $x_{i}(a)\in[m]$, using $k]=\{1,\ldots,k\}$. Let $\mathcal{X}$ be the set of all possible profiles.
@@
[…]
3.1 Embeddings
Traditional feed-forward neural networks require a fixed-size input for learning and inference, corresponding to the size of their input layer (Goodfellow et al. 2016). If we were to learn voting rules using neural networks that take the entire profile as input, then not only does the input layer need to be large $(m\times n)$, but it also prevents scaling up as the number of voters grows without resorting to more complex models (sequential or PINs) which are significantly larger, harder to train, and slower for inference (Anil and Bao 2021). Similarly, if the number of voters shrinks, then the profile would have to be padded carefully to preserve performance. To learn rules that are agnostic to the number of voters we need embeddings of a fixed-size that retain relevant information for profiles with any number of voters. Naturally, different embeddings preserve different information from the original profile, leading to different efficacy when learning different rules and axioms. Note that most rules and axioms in the literature are defined for any positive number of voters, so we would like our learned mechanisms to be similarly agnostic.
Figure 1: Each of the three embeddings derived from a ballot profile. Note that the size of the profile (a) grows in $O(mn)$, while each of our embeddings grows with $O(m^{2})$, which is far smaller when $m<<n$. However, our embeddings do not always preserve all of the information in the original profile.
An embedding $T$ is a function $T:\mathcal{X}\rightarrow\mathcal{X}^{\prime}$ mapping profiles to some codomain $\mathcal{X}^{\prime}$. The embeddings we are concerned with are many-to-one mappings. This means multiple different profiles may have the same embedding, i.e. T(X)=T(X~)T(X)=T($\tilde${X}) for some $X,\tilde{X}\in\mathcal{X}$ where $X\neq\tilde{X}$. In other words, $T$ will not be reversible, and $T(X)$ will not preserve all information about $X$. We denote by $\mathcal{F}^{\prime}$ the set of all probabilistic functions of the form $f^{\prime}:\mathcal{X}^{\prime}\rightarrow\Delta(C)$. Note that while we designate $\mathcal{X}$ to always contain strict orders over candidates, the structure of $\mathcal{X}^{\prime}$ will be different for different embeddings. The following three embeddings are drawn from the voting literature, but are not commonly recognized as embeddings in the machine learning literature.